How many electrons must be removed from a neutral metal sphere so that it acquires a charges of +1nC? .If the radius of the sphere is 0.05m, calculate (1) surface density of the charge on the sphere (2) electric intensity at a point outside and close to the surface of the sphere and (3) electric intensity at a point distance 0.08m from the center of the sphere.

Grade:12

1 Answers

Shailendra Kumar Sharma
6 years ago
1 e has charge of 1.6*10^-19
so to get 1nC charge we need to remove =1X10^-9/( 1.6*10^-19) =6.25X10^9 electrons.
Charge density =q/A =1X10^-9/(4*3.14*.05*.05) =31.85nC/m2